# Data-driven Forecasting of Deep Learning Performance on GPUs

### v1

@@ Abstract.
Deep learning kernels exhibit predictable memory accesses and compute patterns, making GPUs’ parallel architecture well-suited for their execution. Software and runtime systems for GPUs are optimized to better utilize the stream multiprocessors, on-chip cache, and off-chip high-bandwidth memory. As deep learning models and GPUs evolve, access to newer GPUs is often limited, raising questions about the performance of new model architectures on existing GPUs, existing models on new GPUs, and new model architectures on new GPUs. To address these questions, we introduce NeuSight, a framework to predict the performance of various deep learning models, for both training and inference, on unseen GPUs without requiring actual execution. The framework leverages both GPU hardware behavior and software library optimizations to estimate end-to-end performance. Previous work uses regression models (Isaev et al., 2023; Li et al., 2023) that capture linear trends or multilayer perceptrons (Yu et al., 2021) to predict the overall latency of deep learning kernels on GPUs. These approaches suffer from higher error percentages when forecasting performance on unseen models and new GPUs. Instead, NeuSight decomposes the prediction problem into smaller problems, bounding the prediction through fundamental performance laws. NeuSight decomposes a single deep learning kernel prediction into smaller working sets called tiles, which are executed independently on the GPU. Tile-granularity predictions are determined using a machine learning approach and aggregated to estimate end-to-end latency. NeuSight outperforms prior work across various deep learning workloads and the latest GPUs. It reduces the percentage error from 198% and 19.7% to 3.8% in predicting the latency of GPT3 model for training and inference on H100, compared to state-of-the-art prior works (Yu et al., 2021; Li et al., 2023), where both GPT3 and H100 were not used to train the framework.
@@ 1. Introduction
GPUs are the canonical platform to be deployed across any Artificial Intelligence (AI) cluster. There are two main reasons for their widespread adoption for deep learning: (1) the parallel compute engines that mitigate the Von-Neumann overheads of general purpose CPUs and (2) closely knit software libraries that exploit the GPU architecture for various Deep Neural Network (DNN) kernels (NVIDIA, ; NVIDIA, 2017). This trend has been further augmented by the availability of specialized AI servers, aiming to minimize the overheads of distributed execution (Nvidia, ; AMD, ).
@@
Figure 1. Growth of AI models and the compute and memory capacity of GPUs. (NVIDIA, ; AMD, ; Krizhevsky et al., 2012; Simonyan and Zisserman, 2015; Devlin et al., 2019; Radford et al., 2019; Raffel et al., 2023; Brown et al., 2020; Smith et al., 2022)
@@
In this rapidly evolving field of AI, constant iterations are continuously introduced over the years. However, these newer GPUs are limited in access - either due to the long lead time (of upto 52 weeks for H100 (Forbes, 2024)) or their prohibitively high cost. Currently, devising new models and the success of model execution is contingent on the availability of underlying hardware. As shown in Figure 1, recent times (since 2013) have witnessed exponential growth in model size, accompanied by a corresponding increase in the peak performance of GPUs. In summary, the prevailing shift towards the widespread deployment of GPU-centric workloads raises a significant challenge: how can we closely estimate the performance of deep learning models on inaccessible GPUs and upcoming models?
@@
The primary challenge in developing a solution capable of closely estimating GPU performance using only the abstract hardware architecture features lies in the fact that the performance of deep learning models on GPUs is intricately tied to low-level hardware details and software optimizations, such as those provided by DNN kernel libraries like CuDNN and Cutlass. Unlike hardware accelerators (Jouppi et al., 2023; Sharma et al., 2016; Mahajan et al., 2016, 2018; Fowers et al., 2018; Chen et al., 2019), GPUs exhibit complex interplay between micro-architectural interactions and execution characteristics, due L1 and L2 caches, DRAM, memory, I/O configurations and warp scheduling. Consequently, while there are various established works (Olyaiy et al., 2023; Kwon et al., 2019) that can closely estimate the performance of DNN kernels on domain-specific accelerators, they are not suitable for GPU performance forecasting. Nevertheless, previous works have aimed to leverage cycle-accurate simulators to provide detailed execution and performance estimates for GPUs (Bakhoda et al., 2009; Khairy et al., 2020; Avalos Baddouh et al., 2021). However, these simulators need to be manually updated for each change in the GPU architecture. As such, developing a cycle accurate simulator is a time-consuming process and using them to simulate the simplest of the model can take a long time. The most popular GPU simulator (Khairy et al., 2020; Avalos Baddouh et al., 2021) can take up to 18 hours to simulate ResNet-50 with a batch size of 256 (Li et al., 2023). ResNet-50 is relatively smaller compared to new-age complex networks such as the LLMs. Thus, it will take a prohibitively long time to use this simulator to predict the performance of future deep learning models.
@@
In this work, we devise a novel prediction mechanism built on top of a machine learning model, designed with consideration of the tile-based execution strategy on Stream Multiprocessors (SMs) in GPUs, and the compute and memory bounds of DNN kernels. Our approach involves utilizing a multi-layer perceptron to encompass both the hardware features of the GPU and the execution patterns of deep learning. We train our model using data collected from a set of readily available older generation GPUs and existing deep learning kernel information, which is then used to predict performance of newer models, both training and inference, on unseen GPUs.
@@
Thus, we develop NeuSight which digresses from prior works as it leverages the insight that popular GPU libraries decompose deep learning kernels into multiple smaller working sets. These small working sets, or tiles, are then dispatched and executed on the GPU independently across the SMs. Based on this observation we partition the end-to-end latency prediction into regular and more manageable sub-problems, predicting the latency at tile-granularity. However, note that the GPU execution is governed by a few performance laws - such as the performance cannot exceed the peak flops or peak memory bandwidth. We capture the above two aspects in our prediction and only use a machine learning model to determine the utilization of the device based on the DNN kernel and GPU properties. The output of the MLP, i.e., utilization of the device, is subsequently used to predict the tile, operator, and the model level latency. Overall, the goal of this entire framework is to distill down to the properties of GPU execution, use machine learning to capture the runtime non-linear behavior, bound the results by fundamental performance laws, and gather the predictions for end-to-end latency predictions. This approach ensures robustness in unseen scenarios, such as new GPUs and models.
To ensure comprehensive prediction for large models that require multiple GPUs, we extend our approach to predict the latency of distributed execution within a server. To do so, we predict the latency of communication operators and integrate these with the per-device execution latencies determined by NeuSight to calculate the overall distributed performance.
@@ 2. Background
In this section, we dive into the background of GPU architecture and how DNNs are executed on it.
@@ 2.1. GPU Architecture
The availability of parallelism in deep learning workloads and their suitability to GPU-architectures has led to co-optimization across various generations of workloads and GPUs. While the focus in this paper is on NVIDIA GPUs and their terminology, it is important to note most GPUs have similar architectures(AMD, 2024). A GPU is comprised of a collection of identical building blocks known as Streaming Multiprocessors (SMs). These SMs are interconnected with L2 cache through an on-chip interconnection network. The L2 cache, in turn, is connected to memory through off-chip I/O. Each SM features multiple processing units and a small L1 cache. The regularity of the GPU architecture makes it scalable to tens of thousands of processing units, allowing it to effectively exploit parallelism in deep learning kernels.
@@ 2.2. Deep Learning Execution
Per-device execution. Machine learning practitioners use frameworks such as PyTorch and TensorFlow (PyTorch, a; Google, ) to represent and deploy a model as a set of layers comprising their corresponding kernels. The dataflow of the kernels is managed by the framework and dispatched to the GPU. The corresponding device library then executes each kernel and returns the results to the framework. For example, NVIDIA GPUs utilize the CuDNN library for kernel execution, abstracting the underlying hardware complexity from the user. In this paper, similar to prior works, we assume kernels execute sequentially on the device, which is also empirically validated for the mainstream machine learning frameworks (Zhu et al., 2020). Thus, the predicted latency of each kernel is aggregated based on the dataflow of the model that executes on each device to determine the per-GPU execution latency.
Distributed per-server execution. Due to the size of the model, training or inference is often executed in a distributed fashion. Pipeline parallel and tensor model parallel are used to split a model across various devices, whereas data parallel is used to improve throughput (Narayanan et al., 2019; Huang et al., 2019; Nvidia, ; Krizhevsky et al., 2017). Regardless, distributed execution warrants communication across devices such as send/receive for pipeline parallel and collective operations like all-reduce for tensor model parallel and data parallel. In this work, NeuSight supports prediction of latency when a model is distributed across GPUs within a single server. Within this server, GPUs can be connected through NVLink or like a DGX box to support the aforementioned parallelization techniques (NVIDIA, a; Nvidia, ). The type of parallelism, its width, and the schedule is provided by the user as an input to NeuSight.
@@
Figure 2. Prediction error of prior works on BMM operator, reported in percentage error. Out of distribution dimensions and GPUs are highlighted. For these results, we trained Habitat and Li et al. models only up to V100, excluding any A100s, H100 and L4.
@@ 3. Motivation
There is a growing need to predict GPU performance for both existing and upcoming models for a variety of reasons: (a) comparing performance across various devices, (b) utilizing estimates to identify GPUs that meet the performance requirements, and (c) evaluating the feasibility of deploying new model architectures on either existing or new GPUs. There are certain prior works that aim to predict the performance of GPUs by employing data-driven methods, and we empirically validate these works and observe that do not generalize well to new GPUs or new models. These drive the insights for our framework, NeuSight.
For motivation, we focus on a single kernel Batched Matrix Multiplication (BMM), commonly found in DNNs, and present end-to-end comparisons (training and inference) in Section 6. We observe that even predicting performance of a single kernel on a GPU requires a careful consideration of hardware architecture and software optimization. This insight is based on empirical validation of prior works that employ machine learning to determine the operator latency.
@@ 3.1. Predicting Performance of Batched Matrix Multiplication Using Prior Works
Habitat classifies operators into two types: kernel-alike which use similar kernels across different GPU architectures, and kernel-varying which use different kernels across different GPU architectures. For kernel-alike operators (such as vector operators), it executes applications on an existing GPU to measure latency and scales the latency with the ratio of target GPU hardware configuration to that of existing GPU. As such, Habitat requires a GPU in-hand to estimate the latency of these operators. For kernel-varying operators (such as BMM), they train multiple Multi-Layer Perceptron (MLP) models to predict the latency of different operators. Their MLP trains on the following GPU features: memory size, memory bandwidth, number of SM cores, Peak FLOPS, and kernel input and output dimensions as input, and the measured latency as the output.
@@
Figure 2 reports the prediction errors for these two works. For Habitat, we utilized the vanilla MLP model provided by the authors. This model was trained on diverse GPUs predating 2018, including V100, 2080Ti, T4, RTX2070, P100, and P4000, with matrix dimensions of up to 1024 and a batch size less than 128. We aim to understand the applicability of this work on newer GPUs and larger dimensions common in newer models. As Figure 2a shows, Habitat faces challenges in generalizing to matrix multiplications with dimensions larger than 1024, resulting in a percentage error of up to 38%. Additionally, Habitat struggles with unseen GPUs, displaying a percentage error of up to 127% on A100.
For Li et al., we gather BMM data on P4, T4, P100, and V100. We apply linear regression to determine the correlation between the FLOP count of the kernel and latency following the procedures outlined in the paper. For GPUs outside the training set (A100, L4, and H100), we use this linear regression to predict the latency. As Figure 2b shows, the linear regression assumption fails on matrix multiplications with smaller dimensions due to under-utilization of the GPU, when the latency does not scale linearly. This work also faces challenges in generalizing to unseen GPUs because of its simplistic extrapolation method, i.e., using a linear ratio of memory bandwidth to infer the effective performance of unseen GPUs.
One potential solution to reduce the prediction error could be to use larger models for the same problem, i.e., predicting the latency of an operator using a machine learning model, with more layers and features. Thus, as a first attempt, we leverage a larger and more complex model with an increased number of layers to capture a broader set of execution patterns. To test the limitations of relying purely on large machine learning models, we extend Habitat’s prediction model to include more layers in the MLP or alternatively use a transformer architecture from prior work, Prime (Kumar et al., 2022). We train these models using the same method specified in Section 6.1 for the Habitat model. BMM operations with dimensions between 1 and 4096 are used as the test set. Table 1 summarizes the prediction error averaged across in-distribution BMM operations (matrix dimensions less than 1024) and out-of-distribution BMM operations (some dimensions larger than 1024). Despite improved performance with larger models, all the evaluated models still show high percentage error exceeding 70% for out-of-distribution cases.
It is evident from the results that naively using machine learning techniques to predict kernel latency does not suffice, as they fail to capture the continuous performance improvements of GPUs resulting from both software and hardware optimizations. This drives the key insight for NeuSight: unlike prior work, it does not directly predict the latency of kernels using machine learning; instead, it distills the kernel execution into tiles, uses a machine learning model to predict the utilization of the device per tile, and then determines the latency based on performance bounds defined by GPU architecture. NeuSight only relies on architectural details readily available for any new or incoming GPU.

### v2

# Forecasting GPU Performance for Deep Learning Training and Inference

@@ Abstract.
Deep learning kernels exhibit a high level of predictable memory accesses and compute patterns, making GPU’s architecture well-suited for their execution. Moreover, software and runtime system for GPUs further enable optimizations that aim to better utilize the stream multiprocessors, on-chip bandwidth, multiple levels of cache hierarchy, and off-chip high-bandwidth memory. In the context of deep learning, the entire space of models and GPUs is constantly evolving, as newer models emerge with simultaneous upgrades to the device. However, access to newer GPUs is often limited, raising important questions about the performance of new model architectures on existing GPUs, existing models on new GPUs, and new model architectures on new GPUs. To address these questions, we introduce NeuSight, a forecasting framework to predict the performance of a diverse range of deep learning models, for both training and inference, on unseen GPUs, without requiring actual execution of the target model on the target GPU. The framework leverages both GPU hardware behavior and software library optimizations to estimate the end-to-end performance of these models. We observe that prior works in this area (Isaev et al., 2023; Li et al., 2023; Yu et al., 2021) suffer from high absolute error percentages when forecasting performance on unseen models and new GPUs, as they attempt to model the complex task of predicting the latency of a deep learning kernel on a GPU directly using a machine learning approach. Instead, with NeuSight, we decompose the prediction into smaller problems, while bounding the prediction through fundamental performance laws. NeuSight decomposes a single deep learning kernel prediction into smaller working sets called tiles, which are executed independently on the GPU. Tile-granularity predictions are determined using a machine learning approach and aggregated to estimate the end-to-end latency. As such, NeuSight outperforms prior work across a variety of deep learning workloads and the most up-to-date GPUs. It reduces the percentage error from 121.4% and 30.8% to 2.3% in predicting the latency of GPT3 model for training and inference on H100, in comparison to state of the art prior works (Yu et al., 2021; Li et al., 2023), respectively, where GPT3 and H100 were not used to train any framework.
GPU performance forecasting, ML for Systems, Deep Learning, Training and Inference
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30–April 3, 2025; Rotterdam, Netherlands.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsisbn: 979-8-4007-0698-1/25/03doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Machine learningccs: Computer systems organization Parallel architectures
@@ 1. Introduction
GPUs are the canonical platform to be deployed across any Artificial Intelligence (AI) cluster. There are two main reasons for their widespread adoption for deep learning: (1) the parallel compute engines that mitigate the Von-Neumann overheads of general purpose CPUs and (2) closely knit software libraries that exploit the architecture for various Deep Neural Network (DNN) kernels (NVIDIA, 2024c, 2017). This trend has been further augmented by the availability of specialized AI servers, aiming to minimize the overheads of distributed execution (Nvidia, 2024; AMD, 2024).
@@
Figure 1. Growth of AI models and the compute and memory capacity of GPUs. (NVIDIA, 2024d; AMD, 2024; Krizhevsky et al., 2012; Simonyan and Zisserman, 2015; Devlin et al., 2019; Radford et al., 2019; Raffel et al., 2023; Brown et al., 2020; Smith et al., 2022)
@@
In this rapidly evolving field of AI, continuous iterations are introduced. However, newer GPUs are limited in access - either due to the long lead time (of upto 52 weeks for H100 (Forbes, 2024)) or their prohibitively high cost. Currently, devising new models and the success of their execution is contingent on the availability of underlying hardware. As shown in Figure 1, since 2013, we have witnessed an exponential growth in model size, accompanied by a corresponding increase in the peak performance of GPUs. In summary, the prevailing shift towards the widespread deployment of GPU-centric workloads raises a significant challenge: how can we closely estimate the performance of current and upcoming deep learning models on inaccessible GPUs?
@@
The primary challenge in developing a solution capable of closely estimating GPU performance using only the abstract hardware architecture features lies in the fact that the performance of deep learning models on GPUs is intricately tied to low-level hardware details and software optimizations, such as those provided by DNN kernel libraries such as CuDNN and Cutlass. Unlike hardware accelerators (Jouppi et al., 2023; Sharma et al., 2016; Mahajan et al., 2016, 2018; Fowers et al., 2018; Chen et al., 2019), GPUs exhibit complex interplay between micro-architectural interactions and execution characteristics, due L1 and L2 caches, DRAM, memory, I/O configurations, and warp scheduling. Consequently, while there are various established works (Olyaiy et al., 2023; Kwon et al., 2019) that can closely estimate the performance of DNN kernels on domain-specific accelerators, they are not suitable for GPU performance forecasting. Nevertheless, previous works have aimed to leverage cycle-accurate simulators to provide detailed execution and performance estimates for GPUs (Bakhoda et al., 2009; Khairy et al., 2020; Avalos Baddouh et al., 2021). However, these simulators need to be manually updated for each change in the GPU architecture. As such, developing a cycle accurate simulator is a time-consuming process and using them to simulate the simplest of the model can take a long time. The most popular GPU simulator (Khairy et al., 2020; Avalos Baddouh et al., 2021) can take up to 18 hours to simulate ResNet-50 with a batch size of 256 (Li et al., 2023). ResNet-50 is relatively smaller compared to new-age complex networks such as the LLMs. It will take a prohibitively long time to use this simulator to predict the performance of future deep learning models.
@@
In this work, we devise a novel prediction mechanism based on machine learning, designed with the consideration of tile-based execution strategy on Stream Multiprocessors (SMs) in GPUs, and the compute and memory bounds of DNN kernels. We train our model using data collected from a set of readily available older generation GPUs and existing deep learning kernel information, which is then used to predict performance of newer models, both training and inference, on unseen GPUs.
@@
Thus, we develop NeuSight which digresses from prior works as it leverages the insight that popular GPU libraries decompose deep learning kernels into multiple smaller working sets. These small working sets, or tiles, are then dispatched and executed on the GPU independently across the SMs. Based on this observation we partition the end-to-end latency prediction into regular and more manageable sub-problems, predicting the latency at tile-granularity. However, note that the GPU execution is governed by a few performance laws - such as the performance cannot exceed the peak flops or peak memory bandwidth. We capture the above two aspects in our prediction and only use a machine learning model to determine the utilization of the device based on the DNN kernel and GPU properties. The output of the MLP, i.e., utilization of the device, is subsequently used to predict the tile, operator, and the model level latency. Overall, the goal of this entire framework is to distill down to the properties of GPU execution, use machine learning to capture the runtime non-linear behavior, bound the results by fundamental performance laws, and gather the predictions for end-to-end latency predictions. This approach ensures robustness in unseen scenarios, such as new GPUs and models. To ensure prediction for large models that require multiple GPUs, we extend our approach to predict the latency of distributed execution within a server. To do so, we predict the latency of communication operators and integrate these with the per-device execution latencies determined by NeuSight to calculate the overall distributed performance.
@@ 2.1. GPU Architecture
The availability of parallelism in deep learning workloads and their suitability to GPU-architectures has led to co-optimization across various generations of workloads and GPUs. While the terminology used in this paper is of NVIDIA GPUs, we showcase the applicability of NeuSight to both Nvidia and AMD GPUs. A GPU is comprised of a collection of identical building blocks known as Streaming Multiprocessors (SMs). These SMs are connected to a L2 cache through an on-chip interconnection network. The L2 cache, in turn, is connected to memory through off-chip I/O. Each SM features multiple processing units and a small L1 cache. The regularity of the GPU architecture makes it scalable to tens of thousands of processing units, allowing it to effectively exploit parallelism in deep learning kernels.
@@ 2.2. Deep Learning Execution
Per-device execution. Machine learning practitioners use frameworks such as PyTorch and TensorFlow (PyTorch, 2024b; Google, 2024) to represent and deploy a model as a set of layers. The dataflow of the kernels is managed by the framework and dispatched to the GPU. The corresponding device library then executes each kernel and returns the results to the framework. For example, NVIDIA GPUs utilize the CuDNN library for kernel execution, abstracting the underlying hardware complexity from the user. In this paper, similar to prior works, we assume kernels execute sequentially on the device, which is also empirically validated for the mainstream machine learning frameworks (Zhu et al., 2020). Thus, the predicted latency of each kernel is aggregated based on the dataflow of the model that executes on each device to determine the per-GPU execution latency.
Distributed per-server execution. Due to the size of the model, training or inference is often executed in a distributed fashion. Pipeline parallel and tensor model parallel are used to split a model across various devices, whereas data parallel is used to improve throughput (Narayanan et al., 2019; Huang et al., 2019; Smith et al., 2022; Krizhevsky et al., 2017). Regardless, distributed execution warrants communication across devices such as send/receive for pipeline parallel and collective operations like all-reduce for tensor model parallel and data parallel. In this work, NeuSight supports prediction of latency when a model is distributed across GPUs within a single server. Within this server, GPUs can be connected through NVLink or like a DGX box to support the aforementioned parallelization techniques (NVIDIA, 2024; Nvidia, 2024). The type of parallelism, its width, and the schedule is provided by the user as an input to NeuSight.
@@
Figure 2. Prediction error of prior works on BMM operator, reported in percentage error. Out of distribution dimensions and GPUs are highlighted. For these results, we trained Habitat and Li et al. models only up to V100, excluding any A100s, H100, and L4.
@@ 3. Motivation
There is a growing need to predict GPU performance for both existing and upcoming models for a variety of reasons: (a) comparing performance across various devices, (b) utilizing estimates to identify GPUs that meet the performance requirements, and (c) evaluating the feasibility of deploying new model architectures on either existing or new GPUs. We empirically validate prior works that aim to predict the performance of GPUs by employing data-driven methods and observe that do not generalize well to new GPUs or new models. These drive the insights for our framework, NeuSight.
In this section, for motivation, we focus on a single kernel type Batched Matrix Multiplication (BMM), commonly found in DNNs. We present end-to-end comparisons (training and inference) in Section 6. We observe that even predicting performance of a single kernel on a GPU requires a careful consideration of hardware architecture and software optimization. This insight is based on empirical validation of prior works that employ machine learning to determine the operator latency.
@@ 3.1. Predicting Performance of Batched Matrix Multiplication Using Prior Works
Habitat classifies operators into two types: kernel-alike which use similar kernels across different GPU architectures, and kernel-varying which use different kernels across different GPU architectures. For kernel-alike operators (such as vector operators), it executes applications on an existing GPU to measure latency and scales the latency with the ratio of target GPU hardware configuration to that of existing GPU. As such, Habitat requires a GPU in-hand to estimate the latency of these operators. For kernel-varying operators (such as BMM), they train multiple Multi-Layer Perceptron (MLP) models to predict the latency of different operators. The MLP trains on the following GPU features: memory size, memory bandwidth, number of SM cores, Peak FLOPS, and kernel input and output dimensions as input, and the measured latency as the output.
@@
Figure 2 reports the prediction errors for these two works. For Habitat, we utilize the vanilla MLP model provided by the authors. This model was trained on diverse GPUs predating 2018, including V100, 2080Ti, T4, RTX2070, P100, and P4000, with matrix dimensions of up to 1024 and a batch size less than 128. We aim to understand the applicability of this work on newer GPUs and larger dimensions common in newer models. As Figure 2a shows, Habitat faces challenges in generalizing to matrix multiplications with dimensions larger than 1024, resulting in a percentage error of up to 38%. Additionally, Habitat struggles with unseen GPUs, displaying a percentage error of up to 127% on A100.
For Li et al., we gather BMM data on P4, T4, P100, and V100. We use linear regression to correlate the kernel’s FLOP count with latency, following the paper’s outlined procedures. For GPUs outside the training set (A100, L4, and H100), we use this linear regression to predict the latency. As Figure 2b shows, the linear regression assumption fails on matrix multiplications with smaller dimensions due to under-utilization of the GPU, when the latency does not scale linearly. This work also faces challenges in generalizing to unseen GPUs because of its simplistic extrapolation method, i.e., using a linear ratio of memory bandwidth to infer the effective performance of unseen GPUs.
@@ 3.2. Predicting Performance of Batched Matrix Multiplication with Larger Predictors
One potential solution to reduce the prediction error is to use larger models for the same problem, i.e., predicting the latency of an operator using a machine learning model with more layers and features. We extend the study to use a larger and more complex model with an increased number of layers to capture a broader set of execution patterns. To test the limitations of relying purely on large machine learning models, we add more layers to Habitat’s MLP or alternatively use a transformer architecture from prior work, Prime (Kumar et al., 2022). We train these models using the same method specified in Section 6.1. BMM operations with dimensions between 1 and 4096 are used as the test set. Table 1 summarizes the prediction error averaged across in-distribution BMM operations (matrix dimensions less than 1024) and out-of-distribution BMM operations (some dimensions larger than 1024). Despite improved performance with larger models, all the evaluated models still show high percentage error exceeding 70% for out-of-distribution cases.
It is evident from the results that naively using machine learning techniques to predict kernel latency does not suffice, as they fail to capture the continuous performance improvements of GPUs resulting from both software and hardware optimizations. This drives the key insight for NeuSight: unlike prior work, it does not directly predict the latency of kernels using machine learning; instead, it distills the kernel execution into tiles, uses a machine learning model to predict the utilization of the device per tile, and then determines the latency based on performance bounds defined by GPU architecture. NeuSight only relies on architectural details readily available for any new or upcoming GPU.